OpenStudy (anonymous):

Find the Domain and Range of the function: g(x) = 200(2.06)x

OpenStudy (anonymous):

@John_ES

OpenStudy (john_es):

Do yo mean, \[y=200\cdot2.06^x\]?

OpenStudy (anonymous):

Yes.

OpenStudy (john_es):

Domain are all reals, because exponential functions don't have any irregular points. Range is, \[R=(0,+\infty)\]

OpenStudy (anonymous):

Well it's a function so \[f(x) = 200(2.06)^{2}\]But: \[y = 200(2.06)^{2}\]is the same thing.

OpenStudy (john_es):

Yes, but where you put a 2 there should be a x, if not, the function would be a constant function.

OpenStudy (anonymous):

But how would I write the Domain?

OpenStudy (anonymous):

Right, sorry, I was kind of mixed up.

OpenStudy (john_es):

\[D=\mathbb{R}\]

OpenStudy (anonymous):

Isn't it supposed to be written like ? ≤ x ≤ ?

OpenStudy (anonymous):

And the range: ? ≤ y ≤ ?

OpenStudy (anonymous):

So range = 0 ≤ y ≤ ∞

OpenStudy (anonymous):

@John_ES

OpenStudy (john_es):

Range, \[0<y\]

OpenStudy (anonymous):

My lesson writes it in a different way.

OpenStudy (phi):

to do this problem, we start with \[ g(x) = 200(2.06)^x \] for the domain, you start by assuming x goes from -infinity to + infinity then you look to see if you would either (1) divide by 0 (2) take the square root of a negative number. Neither of those things can happen, no matter what value x is, so the domain is \[ - \infty < x < \infty\]

OpenStudy (anonymous):

Oh, thanks, that's how my lesson described it too.

OpenStudy (phi):

to find the range, test a few values of x. If x is a very negative number \[ g( -10000) = 200(2.06)^{-10000} =\frac{200}{2.06^{10000}} \] 200 divided by a very big number is close to 0... notice we never get to zero, but we can get very close. so the range is 0 < y do the same thing for large x and we see y approaches + infinity the total range is 0 < y < infinity

OpenStudy (anonymous):

Ohh, thanks so much.

OpenStudy (john_es):

And what about, \[g(100000000)=200\cdot(2.06)^{100000000}\] Range must be y>0. The function diverges when x grew.